OpenStudy (anonymous):

as a car passes the point A on a straight road its speed is 10 ms the car moves with constant acceleration a ms square along the road for T seconds until it reaches the point b where its speed is V ms the car travels at this speed for further 10 seconds when it reaches the point c from c it travels for a further T seconds with constant acceleration 3ams square until it reaches a speed 20 ms at the point d. given that the distance between A and D is 675 m find the value of a and t. sketch the graph for t,v for the motion and show that v=12.5

OpenStudy (anonymous):

from the kinematics we know that S(ab)=10t+1/2at^2 S(bc)=10(10+at) S(cd)=t(10+at)+3/2t^2 S(ad)=S(ab)+S(bc)+S(cd)=675m (10+at)+3t=20m/s you can solve a and t

OpenStudy (anonymous):

is that right ? :D